# Intrinsic exceptional point – a challenge in quantum theory

### v1

@@ 1 Introduction
They felt motivated by the instability of the IC spectrum under perturbations [2]. They were able to complement such a numerically supported observation by several rigorous mathematical proofs (cf. also [3]). They found that “the eigenvectors of the imaginary cubic oscillator do not form a Riesz basis” [1]. In spite of having spectrum which is real, discrete and bounded below [4, 5, 6], the manifestly non-Hermitian IC Hamiltonian appeared not to be tractable, in the Riesz-basis sense, as diagonalizable.
@@
Siegl with Krejčiřík concluded that “there is no quantum-mechanical Hamiltonian associated with it” [1]. The same authors also recalled the standard mathematical terminology as introduced in Kato’s book [7] and they reformulated their conclusion: “In the language of exceptional points, the imaginary cubic oscillator possesses an ‘intrinsic exceptional point’ ” which is, as a singularity, “much stronger than any exceptional point associated with finite Jordan blocks” [1].
@@
These words are truly challenging, having also motivated our study of the role of IEPs in the deepest conceptual foundations of the contemporary quantum physics. It makes sense to add that Siegl with Krejčiřík only introduced the concept via the above-cited remark, i.e., without giving a formal definition. They only specified IEP as an $N=\infty$ descendant of the standard “exceptional points associated with a finite Jordan block”, i.e., of the conventional exceptional point of order $N$ (EPN). In this sense the linear differential operator $H^{(IC)}$ is really “essentially different with respect to self-adjoint Hamiltonians” [1].
@@
The problem of interpretation of all of the non-standard, IEP-related quantum bound-state problems remains, at present, open. In our contribution to the currently running discussion of this topic (cf. also [8]) we will start, in section 2, from a brief account of what is known about the linear-algebraic EPN analogues of the ordinary differential IEP Eq. (1). We will recall the recent results of the studies of parameter-dependent $N$ by $N$ matrix quantum Hamiltonians $H^{(N)}(g)$ at finite $N<\infty$, and we will emphasize the possibility and relevance of their association with their canonical representation by an $N$ by $N$ matrix Jordan block. Thus, we will consider the toy-model quantum Hamiltonians $H^{(N)}(g)$ admitting a singular limit when $g\to g^{(EPN)}$.
@@
In the latter limit the operator $H^{(N)}(g)$ itself ceases to be diagonalizable and, hence, it ceases to be acceptable as an eligible quantum Hamiltonian. In section 3 we will emphasize that many of its properties become really reminiscent of the IEP features of the differential-operator model (1) where the corresponding Hilbert space of states is infinite-dimensional, $N=\infty$. We will remind the readers of the existing results concerning physical meaning and impact of the EPN-related finite-dimensional models. We will explain that in many (often called “quasi-Hermitian” [9, 10]) quantum models of such a type the limiting transition $g\to g^{(EPN)}$ can be interpreted as one of the most natural realizations of a genuine quantum phase transition (cf., e.g., the description of a class of exactly solvable models of such a process in [11]).
@@
In section 4 the emphasis will be shifted to the $N\to\infty$ scenarios and to the existence of several very useful analogies between both of the IEP and EPN singular extremes. We will point out that in such a comparison the key role of a methodical guide may be expected to be played by (possibly, amended) perturbation theory. Interested readers will be recommended to find a phenomenologically oriented inspiration as well as many related technical details in the older paper [12]. What is important is that the authors studied there a fairly realistic non-Hermitian Hamiltonian describing an N-particle Bose-Einstein condensate generated by a sink and a source in interaction. Using a combination of several complementary numerical as well as analytic and perturbation methods they managed to detect the presence of the EPN singularities in their model. They also revealed and explained that under small perturbations these singularities did unfold in a very specific manner. Their results appeared encouraging because, as they mentioned, the EPN-related problems are truly challenging and their “further investigations …remain tasks for future research.” In our present paper, we just decided to follow the recommendation.
@@
In sections number 5 and 6 we will, in particular, address our main technical task, viz., an IEP-related generalization of the well known perturbation-theory-based description of a generic unitary quantum system near its singularity. We will succeed in showing that many known technical tricks used and tested near EPN at $N<\infty$ can immediately be transferred to the technically more difficult quantum-dynamical scenarios in which the generic Hamiltonians $H(g)$ lie very close to an IEP limiting extreme.
@@
Ultimately, in our last two sections 7 and 8 and also in the series of six brief Appendices we will complement our considerations by several quantum-physics-oriented contextual remarks.
@@ 2 Conventional exceptional points associated with finite Jordan blocks
This makes the IC model important as a genuine methodical as well as conceptual challenge. Here, we intend to propose and advocate the idea that the resolution of the problem could be guided by the EPN-related “good basis” problem and by the existence of parallels between quantum systems near their respective IEP and EPN singularities.
@@ 2.1 The phenomenon of EPN degeneracy
In an opposite direction of argumentation one could recall the existence of exactly solvable quasi-Hermitian $N$ by $N$ matrix models $H^{(N)}(g)$ of paper [11] for which there exists a vicinity of $g^{(EPN)}$ (say, ${\cal D}$) inside which the respective quantum system is found to admit the standard physical probabilistic interpretation. The diagonalizability of the Hamiltonians $H^{(N)}(g)$ plus a few other technical assumptions (concerning the spectrum) then imply that, in principle at least, we may construct all of the bound-state solutions of the so called time-independent Schrödinger equation
@@
  • [1] for all of the “acceptable” $g\neq g^{(EPN)}$ lying in a vicinity of the EPN value, $g\in{\cal D}$, the normalized eigenvectors $|\psi_{n}(g)\rangle$ of $H^{(N)}(g)$ are getting almost parallel to each other.
@@ 3 The mechanism of unfolding of the EPN degeneracy
3.1 Hypothesis of admissibility of the choice of $g\neq{g}^{(EPN)}$
The motivation of the above-outlined amendment of the basis is twofold. Firstly, it enables us to re-read our Schrödinger-like Eq. (6) as an equivalent linear-algebraic relation
@@ 3.1 Hypothesis of admissibility of the choice of $g\neq{g}^{(EPN)}$
i.e., as a definition of a canonical $N$ by $N$ matrix Jordan-block representation of any non-Hermitian Hamiltonian of interest at its EPN singularity. Secondly, the amended basis might find an application in a reformulation of standard perturbation theory. In such a reformulation, the role of the unperturbed Hamiltonian could have been played by its unphysical, singular EPN limit. The trick is that we use the columns of ${R}^{(EPN)}$ as unperturbed basis, assuming only that, in the overall perturbation-theory spirit, the perturbed system acquires a standard phenomenological interpretation for the parameters $g$ lying inside a suitable “physical” vicinity ${\cal D}$ of the EPN singularity.
The latter philosophy is to be advocated and used in what follows. We will only assume the knowledge of the transition matrix ${R}^{(EPN)}$ at an exceptional point of finite order and we will only extend the use of this basis also to some vicinity of the singularity. This will enable us to invert the limiting process $g\to{g}^{(EPN)}$ and to consider the original Hamiltonians at some $g\neq{g}^{(EPN)}$. Our knowledge of transition matrix will yield the model in a simplified form described as a perturbation of a Jordan block matrix,
@@ 3.2 The possibility of keeping the perturbed spectrum real
In papers [17, 18] we considered the above-mentioned quantum-dynamics scenarios and, thoroughly, we studied there the criteria of smallness of the perturbations $V^{(N)}$. We showed that the conditions of the stability and unitarity of the system can be given a mathematically as well as phenomenologically consistent form.
For an illustration of details let us set, conveniently, $E^{(EPN)}=0$ and let us consider the bound-state problem as a perturbation of its EPN limit,
$\left[{J}^{(N)}(0)+\lambda\,V^{(N)}\right]\,|{\Psi}(\lambda)\rangle=\epsilon(\lambda)\,|{\Psi}(\lambda)\rangle\,$ (10)
with the energy levels counted, whenever needed, by a subscript or superscript. We will never use this index, nevertheless, keeping it just dummy. We will rather introduce another subscript which will run, say, from $1$ to $N$ and which will number the components $\Psi_{j}$ of the ket vector $|{\Psi}\rangle$ (here we are also dropping the argument $\lambda$ as redundant).
This convention enables us to fix the norm of $|{\Psi}\rangle$ by the choice of ${\Psi}_{1}=1$ and to define another, “shifted” column vector
@@
is just an inverse of the two-diagonal matrix
@@
Finally we select the first column of the matrix in Eq. (10) and denote it by another dedicated symbol,
@@
where the symbol $Z$ stands just for a modified form of the matrix of perturbation,
@@
In paper [17] we proved that the construction of the perturbation corrections now becomes reduced to the self-consistency condition
@@
In the same reference, interested readers may also find an explicit form of the construction in the leading-order approximation. In fact, its basic aspects are also worth recalling because they immediately help to clarify the meaning of the rather vague above-mentioned assumption of the “sufficient smallness” of the perturbation. In essence, it is sufficient to employ the Taylor-series expansion of the resolvent which yields formula
@@
Such a form of the wave-function-representing ket-vector still depends on the variable parameter $\epsilon$ but, ultimately, all of the eligible values of $\epsilon$ become fixed by constraint (18).
The latter constraint plays the role of secular equation which has, after all, just the scalar vector-component form
@@
In the last step of the construction, obviously, we just have to solve such an explicit transcendental equation in order to get all of the alternative perturbation-generated energy corrections $\epsilon$. In a direct dependence on the model in question, precisely the study of the roots of this equation also offers the criterion of the reality of the whole spectrum in the leading-order approximation.
4 Limit $N\to\infty$ and the anomalous IEP Hamiltonians
@@
4.1 The IEP-related asymptotic degeneracy of eigenvectors
In the analysis of IEP singularities a central role is certainly played by the phenomenon of the asymptotic confluence of infinitely many eigenvectors, not accompanied by the confluence of the eigenvalues. Indeed, in [8] we read that “for matrices approaching an exceptional point it is known [19] that the corresponding eigenvectors are tending to coalesce. For the infinite-dimensional Hilbert space (and Krein space) setup of the IC model, the eigenfunctions of the Hamiltonian having diverging projector norms and asymptotically approaching a PT phase transition region at spectral infinity signal a possible tendency toward collinearity and isotropy of an infinite number of these eigenfunctions” [8].
The EPN - IEP parallels are certainly incomplete so that in the latter case, an amendment of the notation might prove useful. The first mathematical subtlety which we will have to keep in mind is that for a generic IEP model the spectrum itself remains non-degenerate so that, in a way sampled by the IC example, the generic IEP Schrödinger equation
@@ 4.1 The IEP-related asymptotic degeneracy of eigenvectors
can still be considered analogous to its EPN partners. Once the spectrum is found real and discrete (which is precisely the case of the illustrative IC Schrödinger Eq. (1)), the same property also characterizes the formally independent Hermitian conjugate Schrödinger equation problem
@@
at $M\gg 1$. Similarly, the degeneracy concerns also the left alias  “brabra” [20] eigenvectors $\langle\!\langle\psi_{n}^{(IEP)}|$ of the same $H^{(IEP)}$. Often we rather refer to their conjugate form $|\psi_{n}^{(IEP)}\rangle\!\rangle$ of “ketket” eigenvectors of conjugate $\left[H^{(IEP)}\right]^{\dagger}$. In this representation we have
@@
The phenomenon of the confluences (23) and (24) finds its formal predecessor in the finite-dimensional case in which, during the transition to singularity $g\to g^{(EPN)}$ the elements of the $N-$plet of eigenvectors of any preselected $N$ by $N$ Hamiltonian matrix $H=H^{(N)}(g)$ really lose their mutual linear independence. Still, the analogy of a genuine IEP system with the IEP-mimicking $N=\infty$ EPN extreme is incomplete since in the former case the spectrum remains non-degenerate. A more explicit analysis is necessary.
@@ 4.2 Canonical representation of $H^{(IEP)}$
In this arrangement the partitioning of the basis may be characterized by the respective projectors $P$ (on the first $K$ lowest eigenstates) and $Q=I-P$. In a certain parallel with EPN, a key technical challenge will now lie in a weakening or removal of the asymptotic degeneracies (23) and (24).

### v2

@@ 1 Introduction
They felt motivated by the instability of the IC spectrum under perturbations [2]. They were able to complement such a numerically supported observation by several rigorous mathematical proofs (cf. also [3]). They found that “the eigenvectors of the imaginary cubic oscillator do not form a Riesz basis” [1]. In spite of having spectrum which is real, discrete and bounded below [4, 5, 6], the manifestly non-Hermitian IC Hamiltonian appeared not to be, in the Riesz-basis sense, diagonalizable.
@@
Siegl with Krejčiřík concluded that “there is no quantum-mechanical Hamiltonian associated with it” [1]. The same authors also recalled the standard mathematical terminology and they reformulated their conclusion: “In the language of exceptional points, the imaginary cubic oscillator possesses an ‘intrinsic exceptional point’ ” which is, as a singularity, “much stronger than any exceptional point associated with finite Jordan blocks” [1].
@@
These words are truly challenging, having also motivated our study of the role of IEPs in the deepest conceptual foundations of the contemporary quantum physics. It makes sense to add that Siegl with Krejčiřík only introduced the concept via the above-cited remark, i.e., without giving a formal definition. They specified IEP as an $N=\infty$ descendant of the conventional exceptional point of order $N$ (EPN, [7]). In this sense the linear IEP differential operator $H^{(IC)}$ is really “essentially different with respect to self-adjoint Hamiltonians” [1].
The problem of interpretation of all of the non-standard, IEP-related quantum bound-state problems remains, at present, open. In our contribution to the currently running discussion of this topic (cf. also [8]) we will study and describe, more deeply, the parallels as well as differences between the two (viz., the IEP and EPN) concepts.
@@
We will start, in section 2, from a brief account of what is known about the linear-algebraic EPN analogues of the ordinary differential IEP Eq. (1). We will consider a class of Hamiltonians (depending on a real or complex parameter $g$) which admit a singular EPN limit when $g\to g^{(EPN)}$. We will recall a few recent results of the studies of this problem in which a suitable parameter-dependent $N$ by $N$ matrix quantum Hamiltonian $H^{(N)}(g)$ is considered at a finite $N<\infty$. We will emphasize the possibility and relevance of its canonical representation by an $N$ by $N$ matrix Jordan block when $g\to g^{(EPN)}$.
@@
In the latter limit, operator $H^{(N)}(g)$ ceases to be diagonalizable and, hence, it ceases to be acceptable as an eligible quantum Hamiltonian. In section 3 we will emphasize that many of its properties become really reminiscent of the IEP features of the differential-operator model (1) where the corresponding Hilbert space of states is infinite-dimensional, $N=\infty$. We will remind the readers of the existing results concerning physical meaning and impact of the EPN-related finite-dimensional models. We will explain that in many (often called “quasi-Hermitian” [9, 10]) quantum models of such a type the limiting transition $g\to g^{(EPN)}$ can be interpreted as one of the most natural realizations of a genuine quantum phase transition (cf., e.g., the description of a class of exactly solvable models of such a process in [11]).
@@
In section 4 the emphasis will be shifted to the $N\to\infty$ scenarios and to the existence of several very useful analogies between both of the IEP and EPN singular extremes. We will point out that in such a comparison the key role of a methodical guide may be expected to be played by (possibly, amended) perturbation theory. Interested readers will be recommended to find a phenomenologically oriented inspiration as well as many related technical details in older paper [12]. The authors studied there a fairly realistic non-Hermitian Hamiltonian describing an N-particle Bose-Einstein condensate generated by a sink and a source in interaction. Using a combination of several complementary numerical as well as analytic and perturbation methods they managed to detect the presence of the EPN singularities in their model. They also revealed and explained that under small perturbations these singularities did unfold in a very specific manner.
@@
These results appeared encouraging because, as the authors mentioned, the “further investigations” of the EPN-related problems “remain tasks for future research.” In our present paper, we just decided to follow the recommendation. In sections number 5 and 6 we will, in particular, address the main technical challenge and we will propose an IEP-related generalization of the well known perturbation-theory-based description of a generic unitary quantum system near its IEP singularity. We will succeed in showing that many known technical tricks used and tested near EPN at $N<\infty$ can immediately be transferred to the quantum-dynamical scenarios in which the generic Hamiltonian $H(g)$ lies very close to its IEP limiting extreme.
@@
In our last two sections 7 and 8 and also in the series of six brief Appendices we will finally complement our considerations by several quantum-physics-oriented contextual remarks.
@@ 2 Conventional exceptional points associated with finite Jordan blocks
This makes the IC model important as a genuine methodical as well as conceptual challenge. Here, we intend to propose and advocate the idea that the resolution of the problem could be guided by another, EPN-related “good basis” problem and by the existence of parallels between quantum systems near their respective IEP and EPN singularities.
@@ 2.1 The phenomenon of EPN degeneracy
In an opposite direction of argumentation one could recall the existence of exactly solvable quasi-Hermitian $N$ by $N$ matrix models $H^{(N)}(g)$ of paper [11] for which there exists a vicinity of $g^{(EPN)}$ (i.e., say, a suitable compact and simply connected real or complex open domain ${\cal D}$ which does not contain $g^{(EPN)}$ of course) inside which the respective quantum system is found to admit the standard physical probabilistic interpretation. For $g\in{\cal D}$, the diagonalizability of Hamiltonians $H^{(N)}(g)$ then implies that we may construct all of the bound-state solutions of the so called time-independent Schrödinger equation
@@
  • [1] for all of the “acceptable” $g\neq g^{(EPN)}$ lying in the “physical”, unitarity-compatible vicinity of the EPN value, $g\in{\cal D}$, the normalized eigenvectors $|\psi_{n}(g)\rangle$ of $H^{(N)}(g)$ are getting almost parallel to each other.
@@ 3 The mechanism of unfolding of the EPN degeneracy
3.1 The hypothesis of admissibility of at least some $g\approx{g}^{(EPN)}$
The purpose of the above-outlined choice of the basis is twofold. Firstly, it enables us to re-read our Schrödinger-like Eq. (6) as an equivalent linear-algebraic relation
@@ 3.1 The hypothesis of admissibility of at least some $g\approx{g}^{(EPN)}$
i.e., as a definition of a canonical Jordan-block representation of any non-Hermitian Hamiltonian of interest at its EPN singularity. Secondly, the amended basis will find application in a reformulation of standard perturbation theory. In such a reformulation, the role of the unperturbed Hamiltonian will be played by its unphysical, singular EPN limit. The trick is that we use the columns of ${R}^{(EPN)}$ as unperturbed basis. In the overall perturbation-theory spirit, the perturbed system acquires then a standard phenomenological interpretation for the parameters $g$ lying inside a suitable “physical” vicinity ${\cal D}$ of the EPN singularity.
The latter philosophy is to be advocated and used in what follows. We will only assume the knowledge of the transition matrix ${R}^{(EPN)}$ at an exceptional point of finite order and we will then extend the use of this basis to a vicinity of the singularity. This will enable us to invert the limiting process $g\to{g}^{(EPN)}$ and to consider the original Hamiltonians at some $g\neq{g}^{(EPN)}$. Our knowledge of transition matrix will yield the model described as a perturbation of the Jordan block matrix,
@@ 3.2 The possibility of keeping the perturbed spectrum real
In papers [17, 18] we considered the above-mentioned quantum-dynamics scenarios and we studied there the criteria of smallness of the perturbations $V^{(N)}$. We showed that the conditions of the stability and unitarity of the system can be given a mathematically as well as phenomenologically consistent form.
For illustration let us set $E^{(EPN)}=0$ and let us consider the bound-state problem as a perturbation of its EPN limit,
$\left[{J}^{(N)}(0)+\lambda\,V^{(N)}\right]\,|{\Psi}(\lambda)\rangle=\epsilon(\lambda)\,|{\Psi}(\lambda)\rangle\,.$ (10)
With the energy levels counted, whenever needed, by a subscript or superscript, we will never use this index, keeping it just dummy. We will rather introduce another subscript which will run, say, from $1$ to $N$ and which will number the components $\Psi_{j}$ of the ket vector $|{\Psi}\rangle$ (here we are also dropping the argument $\lambda$ as redundant). This convention enables us to fix the norm of $|{\Psi}\rangle$ by the choice of ${\Psi}_{1}=1$ and to define another, “shifted” column vector
@@
is just an inverse of two-diagonal matrix
@@
Finally we select the first column of the matrix in Eq. (10) and we denote it by another dedicated symbol,
@@
where the symbol $Z$ stands for a modified form of the matrix of perturbation,
@@
In paper [17] we proved that the construction of the perturbation corrections now becomes reduced to self-consistency condition
@@
In the same reference, interested readers may also find an explicit form of the construction in the leading-order approximation.
Its basic aspects are worth recalling because they immediately help us to clarify the meaning of the rather vague assumption of the smallness of perturbation. It is sufficient to employ the Taylor-series expansion of the resolvent which yields formula
@@
Such a wave-function-representing ket-vector depends on the variable parameter $\epsilon$ but, ultimately, all of the eligible values of $\epsilon$ become fixed by constraint (18).
The latter constraint plays the role of secular equation which has the single vector-component form
@@
In the last step of the construction we have to solve such an explicit transcendental equation in order to get all of the alternative perturbation-generated energy corrections $\epsilon$. In a direct dependence on the model in question, precisely the study of the roots of this equation also offers the criterion of the reality of the whole spectrum in the leading-order approximation.
4 Large $N$ and anomalous Hamiltonians
@@
In the analysis of both of the EPN and IEP singularities a central role is certainly played by the phenomenon of the asymptotic confluence of finitely or infinitely many eigenvectors, not accompanied by the confluence of the eigenvalues in the IEP case. This can be found confirmed in [8] where we read that “for matrices approaching an exceptional point, it is known [19] that the corresponding eigenvectors are tending to coalesce. For the infinite-dimensional Hilbert space (and Krein space) setup of the IC model, the eigenfunctions of the Hamiltonian having diverging projector norms and asymptotically approaching a PT phase transition region at spectral infinity signal a possible tendency toward collinearity and isotropy of an infinite number of these eigenfunctions” [8].
4.1 The phenomenon of asymptotic degeneracy of eigenvectors
The EPN - IEP parallels are certainly incomplete. Still, in both cases an amendment of the notation might prove useful. Here, we will follow the notation convention which was proposed in our comprehensive review paper [20]. In this spirit, the first mathematical subtlety which we will have to keep in mind is that for a generic IEP model the spectrum itself remains non-degenerate. Still, in a way sampled by the IC example, the generic IEP Schrödinger equation
@@ 4.1 The phenomenon of asymptotic degeneracy of eigenvectors
can be considered analogous to its finite-dimensional EPN-supporting partners.
Once the spectrum is found real and discrete (which is precisely the case of our illustrative IC Schrödinger Eq. (1)), the same property also characterizes the formally independent Hermitian conjugate Schrödinger equation problem
@@
at $M\gg 1$. Similarly, the degeneracy concerns also the left eigenvectors alias  “brabra” eigenvectors $\langle\!\langle\psi_{n}^{(IEP)}|$ of the same non-Hermitian operator $H^{(IEP)}$. Often we rather refer to their conjugate form $|\psi_{n}^{(IEP)}\rangle\!\rangle$ of “ketket” eigenvectors of conjugate $\left[H^{(IEP)}\right]^{\dagger}$. In this representation we encounter an entirely analogous IEP-related confluence of the eigenvectors,
@@
The phenomenon of the confluences (23) and (24) finds its formal predecessor in the finite-dimensional case in which, during the transition to singularity $g\to g^{(EPN)}$, the elements of the $N-$plet of eigenvectors of any preselected $N$ by $N$ Hamiltonian matrix $H=H^{(N)}(g)$ really lose their mutual linear independence. Still, the analogy of a genuine IEP system with the IEP-mimicking $N=\infty$ EPN extreme is incomplete since in the former case the spectrum remains non-degenerate. A more explicit analysis is necessary.
@@ 4.2 Canonical representation of $H^{(IEP)}$
In this arrangement the partitioning of the basis may be characterized by the projectors $P$ (on the first $K$ lowest eigenstates of $H^{(IEP)}$) and $Q$ (such that the unit operator $I$ in Hilbert space can be decomposed as follows, $I=Q+P$).
In a certain parallel with EPN, a key technical step will now be a suitable perturbation-mediated weakening or removal of the asymptotic degeneracies (23) and (24) of the asymptotic eigenstates of $H^{(IEP)}$.